# Phase-Coherent Dynamics in Intervalley Coherent States

### v1

@@ Abstract
Electron transport driven by the phase coherence and interference of quantum many-body wavefunctions is a fascinating phenomenon with potential technological significance. Superconductivity, for example, enables dissipationless transport through macroscopic phase twisting. Similarly, in charge-density waves, once the phase degree of freedom—representing the collective position of electrons relative to the lattice—is depinned, it generates characteristic broadband noise and intriguing AC-DC interference patterns. In this work, we point out a phase-coherent transport phenomena in the intervalley coherent (IVC) state, also known as the bond-ordered or Kekulé distorted state, frequently reported in rhombohedral multilayer graphene. Under a static magnetic field, the IVC state responds with an oscillating orbital magnetization, inducing an AC Hall effect similar to the AC Josephson effect in superconductors. In this analogy, the magnetic field acts as the DC voltage, while the oscillating magnetization acts as the AC Josephson current. We present detailed microscopic calculations for all the parameters of the phase-number free-energy in rhombohedral trilayer graphene, predicting an oscillation frequency of approximately 12 GHz at 0.1 Tesla. We comment on this phase-coherent transport in twisted homobilayer transition metal dichalcogenides, where the IVC state has been theoretically proposed.
@@
Introduction: Magnetism arising from spin and valley ordering in lightly-doped multilayer graphene systems, both with [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] and without moiré potential [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], has become a well-established phenomenon following the discovery of the integer quantum anomalous Hall effect in twisted bilayer graphene [27, 28, 29, 30]. There are three primary types of spin-valley ordered states: spin-polarized ferromagnets, valley-polarized orbital ferromagnets, and the intervalley-coherent (IVC) state. Spin-polarized ferromagnets typically expand their phase boundaries under in-plane magnetic fields, whereas valley-polarized orbital ferromagnets are known for exhibiting a strong anomalous Hall effect. These states, characterized by Ising-like order parameters, also generate magnetic stray fields detectable via nano-SQUID [31, 32, 33, 34]. In contrast, the intervalley-coherent (IVC) state, characterized by its XY-like order parameter, remains elusive and is typically identified by eliminating other symmetry-broken states and through comparison with mean-field theory [14]. Theories suggest that IVC fluctuations could be the microscopic origins of superconductivity [35, 36] in multilayer graphene and homobilayer moiré transition metal dichalcogenides [37]. The IVC state was previously thought to exhibit no prominent transport signatures and can mainly be detected through STM experiments [38, 39, 40, 41]. We note that the difference in the dielectric screening environment can influence the observable magnetic grounstates in STM versus transport experiments [41, 42, 43].
@@
In this study, we report a unique transport phenomenon associated with the intervalley-coherent (IVC) state: an out-of-plane static magnetic field induces an oscillating orbital magnetization, leading to an AC Hall effect, as shown in Fig. 1. This effect, analogous to AC Josephson effect in a superconductor, can be detected using probes sensitive to finite-frequency Hall conductivity $\sigma_{xy}(\omega)$, such as magneto-optical Kerr effect measurements, and may serve as a definitive diagnostic tool for identifying the IVC state.
@@
$2K$ Exciton Condensate: In multilayer graphene with a low electron density of about $10^{-4}$ electrons per carbon atom (typical in most experiments), electrons occupied a tiny fraction of the Brillouin zone, specifically around the opposite corners, which are denoted as valleys $\tau={K,K^{\prime}}$. To make the exciton picture more transparent, we perform a particle-hole transformation in one of the valleys and denote its transformed vacuum by $|\Psi_{vac}\rangle$. Then, the IVC ground state can be written analogously to a BCS superconductor, except that the pairing occurs between electrons and holes in opposite valleys, rather than between electrons with opposite spins and momenta:
@@
Eq. (2) and (3), along with the superfluid kinetic terms:
@@
form the basis of our analysis on unusual transport phenomena in IVC states. They define a phase-number action, $F_{\text{IVC}}$, that applies to any $2\bm{K}$ exciton-condensate with electrons and holes carrying opposite orbital angular momentum:
@@
Note, $\rho$ is the valley stiffness constant associated with the spatially non-uniformity of the IVC phase, and $\chi^{-1}$ is the inverse susceptibility constant associated with the directional anisotropy in valley degree of freedom [51]. In this article, we focus on a uniform solution to the initial-value problem with initial conditions $\phi(0)=N_{\tau}(0)=0$. As shown in Fig. 2, static magnetic field acts like to a chemical potential for the exciton, inducing a nearly linear, coherent phase rotation of the condensate over time. This phase dynamics then subsequently generates an oscillating exciton number $N_{\tau}(t)$, through intervalley tunneling, leading to an oscillating orbital magnetization, $M(t)=g_{v}\mu_{B}N_{\tau}(t)$. Table. I summarizes the parallels between the AC Josephson effect in a superconductor and the analogous phenomenon in our $2K$ exciton condensate. For small $N_{\tau}(t)/\chi$, the solution is straightforward: $\phi(t)=\omega_{B}t$ and $M(t)=M_{0}[1-\cos(3\omega_{B}t)]$ where $M_{0}=g_{v}(\gamma/\hbar\omega_{B})\mu_{B}$ and $\omega_{B}=g_{v}\mu_{B}B/\hbar$. Note the amplitude (frequency) of $M(t)$ decreases (increases) linearly with external field $B$. The minor nonlinear feature in Fig. 2$(a)$ results from the finite $N_{\tau}(t)/\chi$.
@@
AC response Josephoson current Orbital magnetization
Table 1: Comparison of the AC Josephson effect in superconductors and phase-coherent transport in IVC systems.
Rhombohedral trilayer graphene (RTG): We now demonstrate how to compute all the parameters in $F_{IVC}$ for rhombohedral trilayer graphene using a detailed microscopic $K\cdot p$ model.We start with the many body Hamiltonian $\hat{H}=\hat{H}_{b}+\hat{V}$, where
@@
where $\hat{n}_{\bm{q}}=\sum_{\bm{k},\sigma,\tau,s}c_{\tau,\sigma,s,\bm{k}+\bm{q}}^{\dagger}c_{\tau,\sigma,s,\bm{k}}$ and $V_{\bm{q}}=2\pi k_{e}\tanh{(|\bm{q}|d)}/(\epsilon_{r}|\bm{q}|)$ ) is the dual gated screened Coulomb potential. The layer-dependent anisotropy interaction is suppressed in the strong electric-displacement limit and it is not important for discussing IVC transport. Here, $k_{e}=1.44$eVnm is the Coulomb constant, $\epsilon_{r}=15$ is the screening constant, and $d=5$ nm is the distance from the gate to the material. Next, we solve this Hamiltonian using the Hartree-Fock approximation, $\hat{H}\approx\hat{H}_{MF}$ where the matrix elements of $\hat{H}_{MF}$ are determined self-consistently at effectively zero temperature (with negligible temperature for numerical smearing purposes) and at a fixed electron density $n_{e}$. For detail discussion of the self-consistent Hartree-Fock calculations, including the discussion of Fast Fourier Transformation trick to obtain the self-energy, see Ref. [53]. The groundstate thus obtained at $(n_{e},U)=-2.7\times 10^{11}\text{cm}^{-2},15$ meV is a spin-polarized intervalley coherent (IVC) state, consistent with previous mean-field calculations [54, 55] and in agreement with with experimental [14, 15] outcomes through a process of eliminating other possibilities. A linecut of bandstructure for the IVC running wave is showed in Fig. 3. Next, we introduce a Lagrange multiplier into the Hamiltonian $H$:
@@
and solve it again using mean-field theory to obtain a converged density matrix $\hat{\rho}_{\lambda}=\sum_{nk}n_{F}(\epsilon_{nk}^{\lambda})|\psi_{nk}^{\lambda}\rangle\langle\psi_{nk}^{\lambda}|$ as a function of $\lambda$. The resulting bandstructure with $\lambda=0.05$ is shown in Fig. 4$(a)$.
Figure 3: Energy-momentum dispersion of running waves in IVC metal in RTG. The color bar shows the valley mixing angle, with the inset detailing this angle within the Fermi sea of the top valence band. The dashed line marks the Fermi level.
@@
The introduction of $\lambda$ results in a converged solution with a net valley-imbalance, or in exciton language, a net electron-hole imbalance $N_{\tau}=\text{Tr}[|K\rangle\langle K|-|K^{\prime}\rangle\langle K^{\prime}|)\hat{\rho}_{\lambda}]$. This imbalance causes finite orbital magnetization $M$ due to the opposite orbital moments of electrons and holes, and an energy increase $E=\frac{1}{2}\text{Tr}[(H_{b}+H^{MF})\hat{\rho}_{\lambda}]$ due to the depletion of excitons from the condensate, as shown in Fig. 4. Importantly, this state we simulated with finite $\lambda$ represents a many-body state where the energy increase due to the unpaired excitons of the condensate, $N_{\tau}^{2}/(2\chi)$, is counterbalanced by the energy lowering of the orbital magnetization $-MB$. 111 In valley-pseudospin language, this corresponds to an excited state with a slight valley imbalance, where the valley pseudospin order parameter deviates a little bit away from the XY-plane. This scenario represents a stationary solution of Eq. (6) where $\dot{\phi}=0$ and $N_{\tau}=\chi g_{v}\mu_{B}B$. Here $M$ is computed using the so-called modern theory of orbital magnetization [57]:
@@
Discussions: We reveal a unique transport property in the IVC state and demonstrate their experimental observability by computing all parameters of the IVC phase-number free energy $F_{\text{IVC}}$ using a detailed microscopic theory. IVC has been theoretically identified in a wide range of two-dimensional materials. Notably, valley-projected moiré minibands with finite Chern numbers are common in graphene moiré superlattices [62, 63] and twisted homobilayer transition metal dichalcogenides (TMDs) [47]. Metallic IVC states have recently been proposed as the symmetry-broken phases that can mediate pairing glue for superconductivity observed in $5^{\circ}$ twisted WSe2 [37, 64]. The phase-coherent transport discussed for the IVC metal in RTG can be directly generalized to this system, with modified parameters ($\chi^{-1},g_{v},\Gamma$) readily computed using the methods we outlined above. An interesting scenario arises in incompressible IVC states, which have been proposed to be the groundstate at integer fillings of moiré bands in twisted bilayer graphene [63] and potentially also in twisted homobilayer MoTe2 [47, 65, 66], which shows fractional quantum spin Hall effect [67, 68]. If an incompressible ground state has an IVC order and the moiré minibands in opposite valleys have opposite Chern numbers (±$\pm C$), an external magnetic field induces valley imbalance through topological spectral flow: a magnetic flux $\Phi$ leads to exactly $C\Phi/\Phi_{0}$ occupied states per valley flowing adiabatically across the energy gap into unoccupied states where $\Phi_{0}=h/e$ is the flux quantum. States from valleys with opposite Chern numbers flow in opposite spectral directions, i.e. states with $+|C|$ ($-|C|$) flow upward (downward) in energy. This directional spectral flow creates an valley-imbalance for the edge-states, generating finite orbital magnetization and a net edge current. Consequently, the distinction of our phase-coherent AC responses lies in the oscillating Hall current being delocalized throughout the bulk in IVC metal and being localized along the edges in IVC insulator.
@@
The universality class of $F_{\text{IVC}}$ is XY-like, and its topological defects are phase singularities known as merons [69]. Below the Kosterlitz-Thouless (KT) transition temperature $T_{KT}$, or when $\gamma$ is sufficiently strong, merons and antimerons are bound, preventing phase slippage. However, for $T>T_{KT}$, unbound merons cause valley imbalance through phase slippage. This binding is crucial for the observability of our proposed phase-coherent transport phenomena. We have estimated $T_{KT}$ based on a microscopic calculation of IVC domain wall [52] and found $T_{KT}=(\pi/2)\rho\sim 3$K. This temperature is two orders of magnitude higher than the lowest temperatures accessible in many 2D transport experiments [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], making our proposed phenomena experimentally feasible.
Let us end the manuscript highlighting a related phase-coherent transport phenomenon in a Ising|XY|Ising junction configured as K’|IVC|K. Typically, electron transfer between valleys involves large momentum transfers, generated by intervalley phonons or atomic-scale impurity. However, in an IVC state— electrons exist as a coherent superposition of both valleys—there is neither an energy cost nor a momentum conservation constraint stopping the movement of electrons between valleys. Thus, an infinitesimal bias between K and K’ allows current to flow through an IVC state, potentially manifesting as a zero-bias tunneling current peak [70].

### v2

# Momentum-Space AC Josephson Effect and Intervalley Coherence in Multilayer Graphene

@@ Abstract
Electron transport driven by the phase coherence and interference of quantum many-body wavefunctions is a fascinating phenomenon with potential technological significance. Superconductivity, for example, enables dissipationless transport through macroscopic phase twisting. Similarly, in charge-density waves, once the phase degree of freedom—representing the collective position of electrons relative to the lattice—is depinned, it generates characteristic broadband noise and intriguing AC-DC interference patterns. In this work, we point out a phase-coherent dynamics in the intervalley coherent (IVC) state, also known as the bond-ordered or Kekulé distorted state, frequently reported in rhombohedral multilayer graphene. Under a static magnetic field, the IVC state responds with an oscillating intervalley current, which in turn causes oscillating orbital magnetization, thereby inducing a detectable AC Hall effect. This mechanism mirrors the AC Josephson effect observed in superconductors but now happening in momentum space. In this analogy, the static magnetic field acts as the DC voltage, while the oscillating intervalley current assumes the role of the AC Josephson current. We present detailed microscopic calculations for all the parameters of the phase-number free-energy in rhombohedral trilayer graphene, predicting an orbital magnetization oscillation frequency of approximately 12 GHz at 0.1 Tesla. We comment on this phase-coherent dynamics in other 2D materials like twisted homobilayer transition metal dichalcogenides.
@@
Introduction: Magnetism arising from spin and valley ordering in lightly-doped multilayer graphene systems, both with [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] and without moiré potential [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], has become a well-established phenomenon following the discovery of the integer quantum anomalous Hall effect in twisted bilayer graphene [29, 30, 31, 32]. There are three primary types of spin-valley ordered states: spin-polarized ferromagnets, valley-polarized orbital ferromagnets, and the intervalley-coherent (IVC) state. Spin-polarized ferromagnets typically expand their phase boundaries under in-plane magnetic fields, whereas valley-polarized orbital ferromagnets are known for exhibiting a strong anomalous Hall effect. These states, characterized by Ising-like order parameters, also generate magnetic stray fields detectable via nano-SQUID [33, 34, 35, 36]. In contrast, the intervalley-coherent (IVC) state, characterized by its XY-like order parameter, remains elusive and is typically identified by eliminating other symmetry-broken states [12, 13, 14] and through comparison with mean-field theory [37, 38, 39, 40, 41]. Theories suggest that IVC fluctuations could be the microscopic origins of superconductivity [42, 43] in multilayer graphene and homobilayer moiré transition metal dichalcogenides [44]. The IVC state was previously thought to exhibit no prominent transport signatures and can mainly be detected through STM experiments [45, 46, 47, 48, 49]. We note that the difference in the dielectric screening environment can influence the observable magnetic grounstates in STM versus transport experiments [49, 50, 51].
@@
In this study, we report a unique transport phenomenon associated with the IVC state: an out-of-plane static magnetic field induces an oscillating orbital magnetization, leading to an AC Hall effect, as shown in Fig. 1. This effect, analogous to AC Josephson effect in a superconductor, can be detected using probes sensitive to finite-frequency Hall conductivity $\sigma_{xy}(\omega)$, such as magneto-optical Kerr effect measurements, and may serve as a definitive diagnostic tool for identifying the IVC state.
@@
Momentum-Space AC Josephson Effect: In multilayer graphene with a low electron density of about $10^{-4}$ electrons per carbon atom (typical in most experiments), electrons occupied a tiny fraction of the Brillouin zone, specifically around the opposite corners, which are denoted as valleys $\tau={K,K^{\prime}}$. To make the connection to the BCS wavefunction more transparent, we now perform a particle-hole transformation in one of the valleys and denote its transformed vacuum by $|\Psi_{vac}\rangle$. Then, the IVC groundstate can be written analogously to a BCS superconductor, except that the pairing occurs between electrons and holes in opposite valleys, rather than between electrons with opposite spins and momenta:
@@
Eq. (2) and (3), along with the superfluid kinetic terms,
@@
where $\rho$ and $\chi$ are the phase-stiffness constant and susceptibility to create electron-hole imbalance, form the basis of our analysis on unusual transport phenomena in IVC states. They define a phase-number action, $F_{\text{IVC}}$, that applies to any $2\bm{K}$ exciton-condensate with electrons and holes carrying opposite orbital angular momentum:
@@
Eq. (6) and (7) represent the main results of our article, describing phase-coherent dynamics in IVC. We focus on a uniform solution with the initial conditions $\phi(0)=N_{\tau}(0)=0$. As shown in Fig. 2, static magnetic field acts like to a chemical potential for the exciton, inducing a nearly linear, coherent phase rotation of the condensate over time. This phase dynamics then subsequently generates an oscillating exciton number $N_{\tau}(t)$, through intervalley tunneling, leading to an oscillating orbital magnetization, $M(t)=g_{v}\mu_{B}N_{\tau}(t)$. Table. I summarizes the parallels between the AC Josephson effect in a superconductor and the analogous phenomenon in our $2\bm{K}$ exciton condensate. For small $N_{\tau}(t)/\chi$, the solution is straightforward: $\phi(t)=\omega_{B}t$ and $M(t)=M_{0}[1-\cos(3\omega_{B}t)]$ where $M_{0}=g_{v}(\gamma/\hbar\omega_{B})\mu_{B}$ and $\omega_{B}=g_{v}\mu_{B}B/\hbar$. Note the amplitude (frequency) of $M(t)$ decreases (increases) linearly with external field $B$. The minor nonlinear feature in Fig. 2$(a)$ results from the finite $N_{\tau}(t)/\chi$.
@@
AC response Josephoson current Intervalley current
Table 1: Comparison of the AC Josephson effect occurring in real space in superconductors with its analog in momentum space in IVC states. Oscillating intervalley currents lead to oscillating orbital magnetization.
Rhombohedral trilayer graphene (RTG): We now demonstrate how to compute all the parameters in $F_{\text{IVC}}$ for rhombohedral trilayer graphene using a detailed microscopic $K\cdot p$ model. We primarily focus on RTG as it was the first rhombohedral-stacked graphene in which IVC states were experimentally identified [13]. IVC states have also been observed in other layered graphene structures [59, 60, 61]. We start with the many body Hamiltonian $\hat{H}=\hat{H}_{b}+\hat{V}$, where
@@
where $\hat{n}_{\bm{q}}=\sum_{\bm{k},\sigma,\tau,s}c_{\tau,\sigma,s,\bm{k}+\bm{q}}^{\dagger}c_{\tau,\sigma,s,\bm{k}}$ and $V_{\bm{q}}=2\pi k_{e}\tanh{(|\bm{q}|d)}/(\epsilon_{r}|\bm{q}|)$ ) is the dual gated screened Coulomb potential. The layer-dependent anisotropy interaction is suppressed in the strong electric-displacement limit and it is not important for discussing IVC transport. Here, $k_{e}=1.44$eVnm is the Coulomb constant, $\epsilon_{r}=15$ is the screening constant, and $d=5$ nm is the distance from the gate to the material. Next, we solve this Hamiltonian using the Hartree-Fock approximation, $\hat{H}\approx\hat{H}_{MF}$ where the matrix elements of $\hat{H}_{MF}$ are determined self-consistently at effectively zero temperature (with negligible temperature for numerical smearing purposes) and at a fixed electron density $n_{e}$. For detail discussion of the self-consistent Hartree-Fock calculations, including the discussion of Fast Fourier Transformation trick to obtain the self-energy, see Ref. [63]. The groundstate thus obtained at $(n_{e},U)=-2.7\times 10^{11}\text{cm}^{-2},15$ meV is a spin-polarized intervalley coherent (IVC) state, consistent with previous mean-field calculations [39, 40, 41] and in agreement with with experimental [13, 14] outcomes through a process of eliminating other possibilities. In this calculations, we did not perform a particle-hole conjugation in one of the valleys, following the convention in prior works to allow for direct comparison. However, we note that interpreting the IVC state as an exciton condensate is more natural when such a particle-hole transformation is applied in one of the valleys. A linecut of bandstructure for the IVC running wave is showed in Fig. 3. Next, we introduce a Lagrange multiplier into the Hamiltonian $H$:
@@
and solve it again using mean-field theory to obtain a converged density matrix $\hat{\rho}_{\lambda}=\sum_{nk}n_{F}(\epsilon_{nk}^{\lambda})|\psi_{nk}^{\lambda}\rangle\langle\psi_{nk}^{\lambda}|$ as a function of $\lambda$. (Here $\epsilon_{nk}^{\lambda}$ and $|\psi_{nk}^{\lambda}\rangle$ are the mean-field quasiparticle energy and eigenstate). The resulting bandstructure with $\lambda=0.05$ is shown in Fig. 4$(a)$.
Figure 3: Energy-momentum dispersion of running waves in IVC metal in RTG. The color bar shows cosine of the valley mixing angle, with the inset detailing this angle within the Fermi sea of the top valence band. The dashed line marks the Fermi level.
@@
The introduction of $\lambda$ results in a converged solution with a net valley-imbalance, or in exciton language, a net electron-hole imbalance $N_{\tau}=\text{Tr}[|K\rangle\langle K|-|K^{\prime}\rangle\langle K^{\prime}|)\hat{\rho}_{\lambda}]$. This imbalance causes finite orbital magnetization $M$ due to the opposite orbital moments in opposite valleys, and an energy increase $E=\frac{1}{2}\text{Tr}[(H_{b}+H^{MF})\hat{\rho}_{\lambda}]$ due to the depletion of excitons from the condensate, as shown in Fig. 4. Importantly, this state we simulated with finite $\lambda$ represents a many-body state where the energy increase due to the unpaired excitons of the condensate, $N_{\tau}^{2}/(2\chi)$, is counterbalanced by the energy lowering of the orbital magnetization $-MB$. 111 In valley-pseudospin language, this corresponds to an excited state with a slight valley imbalance, where the valley pseudospin order parameter deviates a little bit away from the XY-plane. This scenario represents a stationary solution of Eq. (6) where $\dot{\phi}=0$ and $N_{\tau}=\chi g_{v}\mu_{B}B$. Here $M$ is computed using the so-called modern theory of orbital magnetization [65]:
@@
Discussions: The universality class of $F_{\text{IVC}}$ is XY-like, and its topological defects are phase singularities known as merons [70]. Below the Kosterlitz-Thouless (KT) transition temperature $T_{KT}$, or when $\gamma$ is sufficiently strong, merons and antimerons are bound, preventing phase slippage. However, for $T>T_{KT}$, unbound merons cause valley imbalance through phase slippage. This binding is crucial for the observability of our proposed phase-coherent transport phenomena. We have estimated $T_{KT}$ based on a microscopic calculation of IVC domain wall [62] and found $T_{KT}=(\pi/2)\rho\sim 3$K. This temperature is two orders of magnitude higher than the lowest temperatures accessible in many 2D transport experiments [16, 17, 18, 13, 14, 15, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], making our proposed phenomena experimentally feasible.
The oscillating orbital magnetization in IVC can decay over time due to dissipative tunneling channels. In superconducting junctions, such dissipation arises from quasiparticle tunneling across the barrier, even at voltages below the superconducting gap. In our momentum-space analog of the AC Josephson effect, atomic-scale disorder can play a similar role by inducing random valley transfer, effectively introducing a "shunt resistance" between the $K$ and $K^{\prime}$ valleys. However, atomic-scale resolution STM studies [47, 48, 49, 71] have definitively identified extensive regions within graphene samples displaying IVC order, free from atomic-scale defects that could generate intervalley scattering.
In summary, we uncovered an interesting phase-coherent dynamics in IVC and demonstrate their experimental observability by computing all parameters of the IVC phase-number free energy $F_{\text{IVC}}$ using a detailed microscopic theory. IVC has been theoretically identified in a wide range of two-dimensional materials. Notably, valley-projected moiré minibands with finite Chern numbers are common in graphene moiré superlattices [72, 10] and twisted homobilayer transition metal dichalcogenides (TMDs) [55]. Metallic IVC states have recently been proposed as the symmetry-broken phases that can mediate pairing glue for superconductivity observed in $5^{\circ}$ twisted WSe2 [44, 73]. The phase-coherent transport discussed for the IVC metal in RTG can be directly generalized to this system, with modified parameters ($\chi^{-1},g_{v},\gamma$) readily computed using the methods we outlined above. An interesting scenario arises in incompressible IVC states, which have been proposed to be the groundstate at integer fillings of moiré bands in twisted graphene [10, 11, 12] multilayers and potentially also in twisted homobilayer MoTe2 [55, 74, 75], which shows fractional quantum spin Hall effect [76, 77]. If an incompressible ground state has an IVC order and the moiré minibands in opposite valleys have opposite Chern numbers (±$\pm C$), an external magnetic field induces valley imbalance through topological spectral flow: a magnetic flux $\Phi$ leads to exactly $C\Phi/\Phi_{0}$ occupied states per valley flowing adiabatically across the energy gap into unoccupied states where $\Phi_{0}=h/e$ is the flux quantum. States from valleys with opposite Chern numbers flow in opposite spectral directions, i.e. states with $+|C|$ ($-|C|$) flow upward (downward) in energy. This directional spectral flow creates an valley-imbalance for the edge-states, generating finite orbital magnetization and a net edge current. Consequently, the distinction of our phase-coherent AC responses lies in the oscillating Hall current being delocalized throughout the bulk in IVC metal and being localized along the edges in IVC insulator.
Finally, we note that Ref. [78] has shown that weak localization is enhanced in time-reversal-invariant IVC states. If this enhancement can be reliably distinguished from intervalley scattering effects – also present in the normal metallic states – it could provide a useful diagnostic for detecting IVC order. In contrast, the phase-coherent response we propose arises from interaction-driven three-body Umklapp scattering and is fundamentally distinct from disorder-related effects. As a result, our approach offers a more unambiguous route for identifying IVC.